OpenStudy (jennychan12):

Integral of tan^2x/(x^2+1) ?

OpenStudy (jennychan12):

\[\int\limits \frac{ \tan^2 x }{ x^2+1 }dx\]

hartnn (hartnn):

put x = \(\tan^{-1}y\) dx=... ?

OpenStudy (jennychan12):

dx = 1/(1+x^2)

hartnn (hartnn):

i got stuck actually...i think product rule needs to be use here...

OpenStudy (jennychan12):

product rule? isn't that only for derivatives? :0

OpenStudy (anonymous):

This problem would be a lot nicer if it was an arctan, but it's not right @jennychan12 ?

hartnn (hartnn):

this is not easy, even the uv (product) rule fails here. please verify the question ?

OpenStudy (jennychan12):

no. there's no arctan in the question.

OpenStudy (anonymous):

If there is no arctan in the integral, then integration by parts will only make things worse, gradually with each term, I tried with the first 2 terms before I lost faith in that attempt. So an u-substitution sounds the most reasonable, but I don't see any way how it could be applied.

hartnn (hartnn):

same ^

OpenStudy (anonymous):

The denominator looks a bit like a sec^2(x) or like a hyperbolic function. But oh well.

OpenStudy (jennychan12):

well the original question is If f is an antiderivative of \[\frac{ \tan^2 x }{ x^2+1 }\] such that f(1)=1/2, then f(0)=?